OpenStudy (anonymous):

solve the following sytem of equation -5x-3y=-19 7x-2y=8

hero (hero):

Did you try solving each equation for y?

OpenStudy (anonymous):

i dont underrstand how to do it

OpenStudy (anonymous):

x=2 y=3

OpenStudy (anonymous):

sandman u rock my socks!! tommorrow imma study so i can understand it

OpenStudy (anonymous):

want me to explain it step by step?

OpenStudy (lilg132):

yes i want u to :)

OpenStudy (anonymous):

ok give me a min to type it

OpenStudy (lilg132):

thanks

OpenStudy (anonymous):

ok so you have -5x-3y=-19 7x-2y=8 which is a system of equations so you need to cancel on of the variables out so you can find the other and to do that you multiply through and get one of the term to cancel like so (2)(-5x-3y)=-19 which turns to -10x-6y=-38 (-3)(7x-2x)=8 which turn to -21x+6y=-24 so you ys cancel and you add whats left so you get -31x=-62 solve for x x=2 plug the x back into either equation i jsut use the top normally and solve for y and you get y=3

hero (hero):

y = (-19 + 5x)/-3 y = (8-7x)/-2 Then you can set y = y: -3(8-7x) = -2(-19+5x) -24 + 21x =38 -10x 31x=62 x = 2 y=3

OpenStudy (anonymous):

thanks